OpenStudy (anonymous):

Please help. Medal and fan! The diameter of a softball is 3.8 inches. Estimate the amount of leather used to cover the softball. Round the result to one decimal place.

OpenStudy (anonymous):

You basically have to find the Surface area of the circle.

OpenStudy (anonymous):

what formula do I use?

OpenStudy (anonymous):

Surface area = 4pi r^2

OpenStudy (anonymous):

the radius is a half of a diameter.

OpenStudy (anonymous):

which is 1.9

OpenStudy (anonymous):

so 1.9^2

OpenStudy (anonymous):

Surface area = 28.88pi

OpenStudy (anonymous):

Yep, and you multiply that with 4.

OpenStudy (anonymous):

7.6

OpenStudy (anonymous):

1.9*1.9 = 3.61 3.61* 4 = 14.44 14.44 * 3.14 = 45.3416

OpenStudy (anonymous):

If you round that it will be 45.34.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Np :)